OpenStudy (anonymous):

aunched from the ground a coin vertically upwards at a speed of 15 m / s, at the same instant and from a height of 40m, are projected vertically downwards a stone with a speed of 5 m / s is calculated up to the crossing. Is the currency is rising or falling at that instant? Where is the money lost when the maximum height reached if?

OpenStudy (anonymous):

launched**

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):

i think coin has lost in air

OpenStudy (anonymous):

means?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

ok lets do this for coin \[x_{1}=15t-0.5 (g)(t^2)\] for stone \[x_{2}=40-(5t+0.5(g)t^2\] using g=9.8m/sec^2 set \[x_{1}=x_{2}\] \[15t-4.90t^2=40-5t+4.90t^2\] solve for t ( i hope you are familiar with solving Quadratic equations) t=2 sec (neglect the other root of quadratic equation) using first equation time for coin can be calculated will compare these two to decide whether coin is moving downward or upward . vf=vi-gt but vf=0 t=vi/g=15/9.8=1.533 since \[1.533<2\] so coin is \[falling\] (moving downward at that time. now Where is the money lost when the maximum height ? for maximum height i will use third equation . \[2(-g)(h)=v^2_{f}-v^2_{i}\] vf=0 here \[h=\frac{v^2_{i}}{2g}\] \[h=\frac{15^2}{2(9.8)}=11.479m\] i hope you got that .if need any help let me know .

OpenStudy (anonymous):

@Muskan

OpenStudy (anonymous):

this is projectile numericals. . .

OpenStudy (anonymous):

@theyatin do not mind but i think it is not projectile motion. because it is simple one dimensional problem under gravity. while projectile is 2 dimensional.

OpenStudy (anonymous):

jo formula use karty hain bare ajeeb hain meri book may nhi hain :(( 2(−g)(h)=v2f−v2i like this

OpenStudy (anonymous):

doesn't matters dear sami its projectile at angle 90' still it is categorized under projectile motion. . . everyting thrown up is projeted. . .

OpenStudy (anonymous):

@Muskan this formula is just yjird equation of motion you know this formula \[2as=v^2_{f}-v^2_{i}\] for gravitation acceleration it changes to \[2gs=v^2_{f}-v^2_{i}\] i just used h instead of s and (-g) because moving upward so g will be negative did u get that??

OpenStudy (anonymous):

third*

OpenStudy (anonymous):

quest :D

OpenStudy (anonymous):

?